OpenStudy (anonymous):

Integration question: An aquarium 9 m long, 4 m wide, and 2 m deep is full of water. Find the following: hydrostatic pressure on the bottom of the aquarium, hydrostatic force on the bottom of the aquarium, and hydrostatic force on one end of the aquarium

OpenStudy (anonymous):

@iambatman I'm becoming a fan just for your name

OpenStudy (anonymous):

Pressure varies with depth as \[P(h)=h \rho g\] Where g is acceleration due to gravity and rho is density of the liquid Since the liquid is water, we know that it's density is 1g/cm^3 or \[\rho=\frac{1g}{cm^3}=\frac{10^{-3}Kg}{(10^{-2}m)^3}=\frac{10^{-3}Kg}{10^{-6}m^3}=\frac{1000Kg}{m^3}\] \[g=9.8ms^{-2}\] and h is given as \[h=2m\] Using this u can calculate the pressure Force can be calculated as \[P=\frac{F}{A}\] Where A is the area of the bottom surface of the tank \[\therefore F=P \times A=h \rho g \times A=\rho \times g \times (l \times b \times h)=\rho g V\] V=volume of tank

OpenStudy (anonymous):

Thank you @RunawayGalaxy this is a fun problem unfortunately I do not have time to go over this right now but I will help you out later when I do have time. :) You need to know that force = pressure x area

OpenStudy (anonymous):

Ah nice work @Nishant_Garg :)

OpenStudy (anonymous):

That seems a bit more complicated than I was expecting...

OpenStudy (anonymous):

Really all you have to do is plug in rho h and g in the equation \[P=h \rho g\] \[h=2m\] \[\rho=10^3Kgm^{-3}\] \[g=9.8ms^{-2}\] With this you can calculate the pressure Now \[F=P \times A=P \times l \times b\] The area of the bottom floor of the tank is given by length times breadth, \[l=9m\] \[b=4m\]

OpenStudy (anonymous):

Realized how dumb that last question was. Is the force an integral with P varying with respect to h?

OpenStudy (anonymous):

To find force on one end of the tank, notice that \[F=P \times A\] Now the area of the surface on the side of the tank is given by breadth times height and is constant \[A=b \times h\] But from the bottom of the tank to the top, the pressure is varying with height \[P(h)=h \rho g\] Consider a small height element dh, for which the small pressure is given by \[dP=\rho g.dh\] Then the small force is given by \[dF=b \times h \times \rho \times g \times dh\] \[dF=b \rho g(h.dh)\] Now the height is varying from 0 to 2m in the tank so integrating within the limits 0 to 2 we can find the force \[F=\int\limits_{0}^{F}dF=b \rho g \int\limits_{0}^{2}h.dh\]

OpenStudy (anonymous):

So should that look like \[F = 4 * 10^3 * 9.8 \int\limits_{0}^{2} 2 dh\]

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

would you like help with this

OpenStudy (anonymous):

Yasss

OpenStudy (anonymous):

let me do a quick refresher from physics we know pressure = force / area here the force is due to weight force = mass * gravity constant mass = density * volume volume = Area * height P = (m*g) / A = (rho * Volume* g) / A = (rho * A* height * g) / A = rho * h * g

OpenStudy (anonymous):

Yup. Nishant helped me get the answer to the first one, I'm just trying to answer the force questions

OpenStudy (anonymous):

since P = 1 kg/m^3 (9.81 ) * 2 m = 19.62 N/m^2 (also known as pascals) P = F / A Area* P = F 5*8* 19.62 = 784.8 Newtons force on bottom for the force on the side, I see the next part we have to integrate (catching up here)

OpenStudy (anonymous):

784.8? for the force on the bottom?

OpenStudy (anonymous):

F = Pressure * area total force = integral rho * g *y * (dy * 5)

OpenStudy (anonymous):

Webwork says it's not

OpenStudy (anonymous):

right

OpenStudy (anonymous):

are you given special values for gravity

OpenStudy (anonymous):

No, I just assumed it would be 9.8

OpenStudy (anonymous):

I see, I used different values for the dimensions of the aquarium should be the dimensions are 9 m long, 4 m wide, and 2 m Force on bottom = P*A F = rho * g * height * Area of base F = 1 * 9.81 * 2 * (9 * 4) = 706.32

OpenStudy (anonymous):

see if that is correct

OpenStudy (anonymous):

Rho is 1? It still says it's wrong

OpenStudy (anonymous):

No, I'm surprised this question would use any of this, this is from a calc class, not physics

OpenStudy (anonymous):

oh woops, i thought it was 1 kg /m^3

OpenStudy (anonymous):

Thereee we go

OpenStudy (anonymous):

That one worked

OpenStudy (anonymous):

which one worked?

OpenStudy (anonymous):

705600

OpenStudy (anonymous):

ok great

OpenStudy (anonymous):

so for the side I would use the same thing?

OpenStudy (anonymous):

they did F = 1000 * 9.8 * 2 * (9 * 4) = 705600

OpenStudy (anonymous):

for the side, we are asking for the total force on the side of the window, so to speak, of the aquarium

OpenStudy (anonymous):

since pressure acts the same in all directions, it is going to act on the side of the aquarium the same as it acts on the floor, but it changes by depth